# Training Machine Learning models at the Edge: A Survey

### 05 Mar 2024 v1

@@ I Introduction
In the past years, the fields of Artificial Intelligence (AI) Machine Learning (ML) has witnessed significant growth and advancements, demonstrating remarkable success across various industrial applications [1]. The essence of ML has always been shaped by the interplay between data and algorithmic models. Large quantities of data being often required to effectively train ML models. And traditionally, the datasets were collected in Cloud Storage Buckets, Large Databases and Data Lakes, and processed in central cloud servers, to train various machine learning and deep learning models.
Conversely, The rapid proliferation of smart devices and sensors in recent years has led to an explosion of data generation at the edge of the network. With edge devices generating vast quantities of data outside the traditional cloud servers. The raising awareness of the privacy and security concerns that come with sending these data to the cloud, as well the desire to optimize the bandwidth consumption on the increasing number of edge devices and reduce the computational load on cloud servers, have precipitated a paradigm shift towards Edge Computing (EC). In this context, computational processes are decentralized and migrated to edge devices, setting the stage for a novel intersection between machine learning and edge computing.
@@ III Comparison with existing surveys
✓: Paper covers the topic fully for Edge Learning; $\circ$ : Paper covers the topic partially for Edge Learning; •: Paper covers the topic with a focus on both Edge Learning and Inference; \: Paper covers the topic only from an Edge Inference point of view; ✗: Paper does not cover the topic;
@@ IV Requirements for edge learning
    Low Energy Consumption: Energy consumption for edge devices is a crucial consideration, especially in mobile edge computing, considering the limited energy available at such devices. For this reason, ML models trained at the edge, need to be energy efficient, allowing for better computing performance and battery lifetime while ensuring the successful training of the models. Energy efficiency refers to the ability to use less energy to perform a specific task or function, it involves reducing energy waste and optimizing energy consumption.
@@
    Labelled Data Independency: The majority of the data produced at the edge are unlabelled data [32], for that reason, it might be beneficial to use in EL, ML techniques that are fully or partially trained using unlabelled data such as unsupervised learning (Section VI-A), self-supervised learning (Section VI-D) or semi-supervised learning (Section VI-C).
@@
    Task Specific Metrics and Performance: As Edge learning encompass different ML tasks and use-cases, specific metrics and benchmarks are used in general for the evaluation of the performances of the model and are vital for assessing its effectiveness in achieving its intended goals.
@@ V Overview of Edge Learning techniques
Ultimately, the possibility of training an ML model on the edge depends on the resources requirements of the model, and the resources available at the device. With edge devices having increasingly more processing power, energy storage, and memory capacity available [33], it is currently possible to train a small ML model in an edge device without having to particularly optimize it for that regard or distribute the training across multiple devices. For examples KMeans in [34], Self-Organizing Map in [35] and SVM in [36]. However the training of more complex models that require heavier resources, such as neural networks is more challenging at the edge. Therefore, in this section, we will present an overview of techniques used to train more complex ML models in the edge, either through distributing the training to multiple devices, by training part of the model on the cloud and fine-tuning it on the local device or by optimizing or compressing the model to facilitate the training on the edge.
@@
collaborative learning (V-A), text depth=3ex, text width=27ex, minimum height=1cm [Federated Learning (V-A1)] [Split Learning (V-A2)] [Swarm Learning (V-A3)] [Gossip Learning (V-A3)] ] [Adaptive and Fine-tuning
based Technique (V-B), text depth=3ex, text width=27ex, minimum height=1cm) [Transfer Learning (V-B1)] [Incremental Learning (V-B2)] [Meta Learning (V-B3)] ] [Model compression (V-C), text width=27ex, minimum height=1cm [Quantization (V-C2)] [Knowledge Distillation (V-C1)] [Model Pruning (V-C3)] ] [Other optimization
methods (V-D), text depth=3ex, text width=27ex, minimum height=1cm [Binary Neural Networks (V-D1)] [Spiking Neural Networks (V-D2)] [Forward-Forward Algorithm (V-D3)] ] ]
@@ V-A1 Federated Learning
Federated learning (FL), offers a transformative approach to decentralized model training. In the context of edge learning, where data is distributed across numerous edge devices, FL enables collaborative training without centralizing sensitive data [37]. This technique involves training a shared model across these devices by iteratively updating it based on local data, while preserving data privacy and minimizing communication overhead. FL is used extensively for training ML models at the edge [9], and has been adopted for multiple use-cases that ranges from detecting cyberattacks [38, 39] to spam detection [40, 41], or used in smart cities [42, 43] and autonomous vehicles [44]. Overall, privacy preservation is usually the main objective for FL [45, 46, 47].
@@ V-A3 Other Collaborative Learning methods
Several distributed learning techniques have been proposed as alternatives to FL and SL. Swarm learning, an innovative approach integrating artificial and biological intelligence, addresses challenges in distributed ML for the edge. This method efficiently utilizes signal processing and communication techniques to operate in real-time within large-scale edge IoT environments, offering advantages in overcoming communication bottlenecks, diverse data, non-convex optimization, and privacy concerns [113]. As another approach to Swarm learning, CB-DSL [114], a Communication-efficient and Byzantine-robust Distributed Swarm Learning technique, was introduced to deal with non-i.i.d. data issues and Byzantine Attacks. Another noteworthy distributed learning method is Gossip Learning, which, like other collaborative methods, doesn’t require transferring data outside edge devices. However, unlike FL and other methods, Gossip Learning operates without a central server for model aggregation and lacks reliance on central control [115]. Notable extensions to Gossip Learning, such as the one proposed by [116], enhance the algorithm by incorporating additional memory for storing local caches of model updates, making it more suitable for mobile devices.
@@ V-B1 Transfer Learning
In the context of edge learning, TL is a prominent technique used to fine-tune ML models based on local data in an edge device, and can serves as a fully on-device alternative to collaborative learning methods such as FL or SL that distribute the training across different devices [31], here are some state-of-the-art methods for transfer learning techniques for edge learning. TinyTL (Tiny-Transfer-Learning) [30] aims to achieve on-device learning memory efficiency, which is especially important for low memory edge devices. It achieves that by freezing the weights of the model and only learning a memory-efficient bias module, thus removing the need to store the intermediate activations. Similarly, RepNet [118] proposes an intermediate feature re-programming of a pre-trained model with a tiny Reprogramming Network to develop memory-efficient on-device transfer learning, and MobileTL [119] proposes a memory and computationally efficient on-device transfer learning method for models built with Inverted Residual Blocks. [120] propose an edge CNN framework for 5G industrial edge networks, with the CNN model trained in advance in an edge server, which is further fine-tuned based on the limited datasets uploaded from the devices with the aid of Transfer Learning. And [121] proposes a runtime convergence monitor to achieve massive computational savings in the practical on-device training workloads. Multiple approaches also focus on combining TL and FL, to create federated transfer learning algorithms [122], that aim to leverage FL for privacy preservation, and use transfer learning to train a well-performing local model despite users usually having not enough data for that by training the base model with a public dataset and passing it to the federated users to be fine-tuned for the target task [97, 123]. Finally, [124] proposes freeze and reconfigure, a TL method for on-device training of a BERT model.
@@ V-C2 Quantization
Although quantization techniques are usually used to optimize machine learning models before deployment for edge inference [164, 121], it can be used in edge learning in order to simplify the fine-tuning of large models in the edge [31, 165]. Similarly to KD, quantization is often alongside other techniques such as federated learning [166], transfer Learning [31], incremental learning [167] or with other types of techniques [168]. Quantization is used with FL particularly extensively, from one-bit quantization [169, 170, 171] to hierarchical FL [100] and other FL based approaches such as [172, 173, 174, 175, 176, 99]. Other Quantization based methods for training ML models in the edge include, Quantization-Aware Scaling (QAS), a method proposed in [31], to automatically scale the gradient of tensors with different bit-precisions without requiring any fine-tuning, the method was used alongside the Tiny Training Engine and Sparse Update introduced in the same paper [31]. In [177], investigations were made on on-device few shot learning for audio classification, and it was found that quantization helps further in reducing the resource requirements for the training. In the optimizer. Holmes [178], quantization is used in an ingenious way to improve the accuracy by effectively combining different types of quantization (a limitation on the number of quantization bits, fixed-point number, and logarithmic quantization) techniques.
@@ V-D3 Forward-Forward Algorithm
The backpropagation algorithm is essential for training neural networks, but recent studies have proposed alternatives to the algorithm when the available resources are limited, such as the Forward-Forward (FF) [205] algorithm that aims to replace the forward and backward passes of backpropagation by two forward passes that operate in the same way as each other on different data and with opposite objectives. A positive pass operates on real data and adjusts the weights to increase the goodness in every hidden layer, while a negative pass operates on ”negative data” and adjusts the weights to decrease the goodness in every hidden layer [205]. Some other studies have tried to adapt the Forward-Forward algorithm to edge devices, such as µ-FF [206] that propose an on-device variation of the original FF on microcontroller units which tackles the training process with a multivariate Ridge regression approach and allows finding closed-form solution by using the Mean Squared Error (MSE) as loss function. And finally, [207] investigates the improvements in terms of complexity and memory usage brought by PEPITA [208] and the Forward-Forward Algorithm, the results show that the Forward-Forward algorithm reduce memory consumption by 40% on average, however it involves additional computation at inference that, can be costly on microcontrollers.
@@ V-E1 Comparison of the usage of the different techniques
We will first start by making a comparison of the different EL techniques’ over the years, by analysing their academic contributions. The figure 4 present the number of papers for each technique, while the figure 5 references the number of papers (in a logarithmic scale) per technique per year. In order to get the necessary data, We used the advanced search feature of Scopus  [^1] Scopus : https://www.scopus.com/ where we searched the terms ”edge learning”, ”training/learning on the edge/mobile devices”, ”on-device training/learning” and ”on-device adaptation” as well as the relevant keywords for each technique (”Federated Learning”, ”Split Learning”, etc), in the Title, Keywords and Abstract, we made sure to exclude the surveys, conference reviews, books and notes as we are only interested in approaches to train ML models in the edge using the aforementioned techniques. For each family of techniques, we manually reviewed and removed papers that either didn’t provide a contribution, or weren’t about training ML models on edge devices with said technique, despite checking the keywords. Finally, we added a few papers we found manually and either were not available in Scopus or didn’t check the keywords but were relevant in the contest of the analysis. The final number of papers associated with the analysis was 511 papers, and some of these papers were briefly covered in V, note that multiple techniques can be used in a single paper, therefore the total count of techniques in the figures 4 and 5 will exceed 511.
@@
Figure 5: Trend of technique used to train ML models in the edge over tthe years
@@ V-E2 Comparison based on the requirements and needs of EL
As discussed in Section IV, several requirements must be met for edge learning, which can function as imprecise measures for assessing the viability of the families of techniques covered in earlier sections. After reevaluating the prerequisites, we eliminated ”Labelled Data Independence” from the comparison, because it is more relevant to the type of machine learning (ML) employed and the availability of an autolabeling process than to the approaches under consideration. Additionally, we do not consider ”Task-specific metrics and performance” since that covers multiple metrics used in specific tasks; nevertheless, we introduce a ”High Performance” measure to estimate roughly if the strategies positively or negatively affect the performances. For instance, model compression techniques often lower model performance, whereas incremental and meta-learning typically boost it. Consequently, we employ six distinct measures: ”Computational Efficiency,” ”Memory Footprint,” ”Low Energy Consumption,” ”Quick Training Time,” ”Optimized Bandwidth,” and ”High Performance.” Table II compares the various families of techniques against these requirements.
@@ VI-A Unsupervised Learning
The figure 7 represent the main different types of unsupervised learning approaches used for training ML models in the edge, (a) using unsupervised learning algorithms directly on-device with non-collaborative methods [216, 35]; (b) using unsupervised learning methods to assist in the training of collaborative learning approaches, such as clustered federated learning [79, 217, 218] ; and (c) training an unsupervised learning model on the edge in a collaborative [219].
@@
Having an unsupervised learning trained on a single edge device is possible if the edge device has enough computation power and/or the learning algorithm is lightweight and can be trained with low resources, the training with such models is no different from the training on the cloud or other devices, the only difference being the constraint of low resources and data available [35]. Examples of unsupervised learning algorithms trained with this approach include [34] that investigate the application of K-Means on mainstream controllers, and [220] that presents the first dedicated CycleGAN accelerator for energy-constrained mobile applications, achieving a higher throughput-to-area ratio and higher energy efficiency than a GPU. In [216], an unsupervised segmentation was proposed that can be executed on edge devices without the need of annotated data. While [221], proposes TMNet, an approach to solve unsupervised video object segmentation problem at the edge. Finally, [222] propose an FPGA based architecture for a Self-Organization Neural Network capable of performing unsupervised learning on input features from a CNN by dynamically growing neurons and connections in order to perform class-incremental lifelong learning for object classification in the edge.
@@
Collaborative learning approaches like FL are promising solution for training ML models in the edge, However FL the most popular technique on the edge, face multiple challenges such as non-independent and identically collaborative (non-IID) data and uneven computing power [110] and yields suboptimal results when the local clients’ data distributions diverge [79]. To solve these issues, the usage of clustering alongside FL have been proposed multiple times to cluster edge servers with similar environmental data distributions [223]. Among these approaches, Clustered Federated Learning (CFL), introduced by [79] is widely adopted, and it works by exploiting geometric properties of the FL loss surface and grouping the client population into clusters with jointly trainable data distributions. Approaches that build upon CFL have also been proposed, and they include [217, 110, 218, 224]. Other approaches that differ from CFL [79] include [68] that propose a hierarchical over-the-air FL approach which utilizes intermediary servers to form clusters near mobile users; and HPFL-CN [223], a communication-efficient hierarchical personalized federated edge learning framework, that achieves this by utilizing complex network feature clustering to group together edge servers that exhibit similar environmental data distributions. Subsequently, personalized models are trained for each cluster using a hierarchical architecture, resulting in enhanced efficiency. More specifically, HPFL-CN incorporates Privacy-preserving Feature Clustering (PFC) to derive privacy-preserving low-dimensional feature representations for each edge server. This is achieved by mapping the environmental data onto various complex network domains, thereby accurately clustering edge servers with similar characteristics. And Finally, [57] uses unsupervised learning methods on the edge to distinguish between classes across different users, when new classes with completely unseen data distributions emerge on devices in a federated learning setting for audio classification.

### 13 Mar 2024 v2

@@ I Introduction
In the past years, the fields of Artificial Intelligence (AI) and Machine Learning (ML) has witnessed significant growth and advancements, demonstrating remarkable success across various industrial applications [1]. The essence of ML has always been shaped by the interplay between data and algorithmic models. Large quantities of data being often required to effectively train ML models. And traditionally, the datasets were collected in cloud storage buckets, large databases and data lakes, and processed in central cloud servers, to train various machine learning and deep learning models.
Conversely, The rapid proliferation of smart devices and sensors in recent years has led to an explosion of data generation at the edge of the network. With edge devices generating vast quantities of data outside the traditional cloud servers, the raising awareness of the privacy and security concerns that come with sending these data to the cloud, as well the desire to optimize the bandwidth consumption on the increasing number of edge devices and reduce the computational load on cloud servers, have precipitated a paradigm shift towards Edge Computing (EC). In this context, computational processes are decentralized and migrated to edge devices, setting the stage for a novel intersection between machine learning and edge computing.
@@ III Comparison with existing surveys
✓: Paper covers the topic fully for Edge Learning;
$\circ$ : Paper covers the topic partially for Edge Learning;
•: Paper covers the topic with a focus on both Edge Learning and Inference;
\: Paper covers the topic only from an Edge Inference point of view;
✗: Paper does not cover the topic;
@@ IV Requirements for edge learning
    Low Energy Consumption: Energy consumption for edge devices is a crucial consideration, especially in mobile edge computing, considering the limited energy available on such devices. For this reason, ML models trained at the edge, need to be energy efficient, allowing for better computing performance and battery lifetime while ensuring the successful training of the models. Energy efficiency refers to the ability to use less energy to perform a specific task or function. It involves reducing energy waste and optimizing energy consumption.
@@
    Labelled Data Independency: The majority of the data produced at the edge are unlabelled data [32]. For that reason, it might be beneficial to use in EL, ML techniques that are fully or partially trained using unlabelled data such as unsupervised learning (Section VI-A), self-supervised learning (Section VI-D) or semi-supervised learning (Section VI-C).
@@
    Task Specific Metrics and Performance: As Edge learning encompass different ML tasks and use-cases. Specific metrics and benchmarks are used in general for the evaluation of the performances of the model and are vital for assessing its effectiveness in achieving its intended goals.
@@ V Overview of Edge Learning techniques
Ultimately, the possibility of training an ML model on the edge depends on the resource requirements of the model, and the resources available on the device. With edge devices having increasingly more processing power, energy storage, and memory capacity available [33], it is currently possible to train a small ML model in an edge device without having to particularly optimize it for that regard or distribute the training across multiple devices. For example, KMeans in [34], Self-Organizing Map in [35] and SVM in [36]. However the training of more complex models that require heavier resources, such as neural networks is more challenging at the edge. Therefore, in this section, we will present an overview of techniques used to train more complex ML models in the edge, either through distributing the training to multiple devices, by training part of the model on the cloud and fine-tuning it on the local device or by optimizing or compressing the model to facilitate the training on the edge.
@@
collaborative learning (V-A), text depth=2ex, text width=27ex, minimum height=1cm [Federated Learning (V-A1)] [Split Learning (V-A2)] [Swarm Learning (V-A3)] [Gossip Learning (V-A3)] ] [Adaptive and fine-tuning
based technique (V-B), text depth=2ex, text width=27ex, minimum height=1cm) [Transfer Learning (V-B1)] [Incremental Learning (V-B2)] [Meta Learning (V-B3)] ] [Model compression (V-C), text width=27ex, minimum height=1cm [Quantization (V-C2)] [Knowledge Distillation (V-C1)] [Model Pruning (V-C3)] ] [Other optimization
methods (V-D), text depth=2ex, text width=27ex, minimum height=1cm [Binary Neural Networks (V-D1)] [Spiking Neural Networks (V-D2)] [Forward-Forward Algorithm (V-D3)] ] ]
@@ V-A1 Federated Learning
Federated learning (FL), offers a transformative approach to decentralized model training. In the context of edge learning, where data is distributed across numerous edge devices, FL enables collaborative training without centralizing sensitive data [37]. This technique involves training a shared model across these devices by iteratively updating it based on local data, while preserving data privacy and minimizing communication overhead. FL is used extensively for training ML models at the edge [9], and has been adopted for multiple use-cases that range from detecting cyberattacks [38, 39] to spam detection [40, 41], or used in smart cities [42, 43] and autonomous vehicles [44]. Overall, privacy preservation is usually the main objective for FL [45, 46, 47].
@@ V-A3 Other Collaborative Learning methods
Several distributed learning techniques have been proposed as alternatives to FL and SL. Swarm learning, an innovative approach integrating artificial and biological intelligence, addresses challenges in distributed ML for the edge. This method efficiently utilizes signal processing and communication techniques to operate in real-time within large-scale edge IoT environments, offering advantages in overcoming communication bottlenecks, diverse data, non-convex optimization, and privacy concerns [113]. Another approach to swarm learning, CB-DSL [114], a Communication-efficient and Byzantine-robust Distributed Swarm Learning technique, was introduced to deal with Non-IID data issues and byzantine attacks. Another noteworthy distributed learning method is gossip learning, which, like other collaborative methods, doesn’t require transferring data outside edge devices. However, unlike FL and other methods, gossip learning operates without a central server for model aggregation and lacks reliance on central control [115]. Notable extensions to gossip learning, such as the one proposed by [116], enhance the algorithm by incorporating additional memory for storing local caches of model updates, making it more suitable for mobile devices.
@@ V-B1 Transfer Learning
In the context of edge learning, TL is a prominent technique used to fine-tune ML models based on local data in an edge device, and can serve as a fully on-device alternative to collaborative learning methods such as FL or SL that distribute the training across different devices [31]. Here are some state-of-the-art methods for transfer learning techniques for edge learning. TinyTL (Tiny-Transfer-Learning) [30] aims to achieve on-device learning memory efficiency, which is especially important for low memory edge devices. It achieves that by freezing the weights of the model and only learning a memory-efficient bias module, thus removing the need to store the intermediate activations. Similarly, RepNet [118] proposes an intermediate feature re-programming of a pre-trained model with a tiny Reprogramming Network to develop memory-efficient on-device transfer learning, and MobileTL [119] proposes a memory and computationally efficient on-device transfer learning method for models built with Inverted Residual Blocks. [120] propose an edge CNN framework for 5G industrial edge networks, with the CNN model trained in advance in an edge server, which is further fine-tuned based on the limited datasets uploaded from the devices with the aid of transfer learning. And [121] proposes a runtime convergence monitor to achieve massive computational savings in the practical on-device training workloads. Multiple approaches also focus on combining TL and FL, to create federated transfer learning algorithms [122], that aim to leverage FL for privacy preservation, and use transfer learning to train a well-performing local model despite users usually having not enough data for that by training the base model with a public dataset and passing it to the federated users to be fine-tuned for the target task [97, 123]. Finally, [124] proposes freeze and reconfigure, a TL method for on-device training of a BERT model.
@@ V-C2 Quantization
Although quantization techniques are usually used to optimize machine learning models before deployment for edge inference [164, 121], it can be used in edge learning in order to simplify the fine-tuning of large models in the edge [31, 165]. Similarly to KD, quantization is often alongside other techniques such as federated learning [166], transfer Learning [31], incremental learning [167] or with other types of techniques [168]. Quantization is used with FL particularly extensively, from one-bit quantization [169, 170, 171] to hierarchical FL [100] and other FL based approaches such as [172, 173, 174, 175, 176, 99]. Other Quantization based methods for training ML models in the edge include, Quantization-Aware Scaling (QAS), a method proposed in [31], to automatically scale the gradient of tensors with different bit-precisions without requiring any fine-tuning, the method was used alongside the Tiny Training Engine and Sparse Update introduced in the same paper [31]. In [177], investigations were made on on-device few shot learning for audio classification, and it was found that quantization helps further in reducing the resource requirements for the training. In the Holmes optimizer [178], quantization is used in an ingenious way to improve the accuracy by effectively combining different types of quantization (a limitation on the number of quantization bits, fixed-point number, and logarithmic quantization) techniques.
@@ V-D3 Forward-Forward Algorithm
The backpropagation algorithm is essential for training neural networks, but recent studies have proposed alternatives to the algorithm when the available resources are limited, such as the Forward-Forward (FF) [205] algorithm that aims to replace the forward and backward passes of backpropagation by two forward passes that operate in the same way as each other on different data and with opposite objectives. A positive pass operates on real data and adjusts the weights to increase the goodness in every hidden layer, while a negative pass operates on ”negative data” and adjusts the weights to decrease the goodness in every hidden layer [205]. Some other studies have tried to adapt the Forward-Forward algorithm to edge devices, such as µ-FF [206] that propose an on-device variation of the original FF on microcontroller units which tackles the training process with a multivariate Ridge regression approach and allows finding closed-form solution by using the Mean Squared Error (MSE) as loss function. And finally, [207] investigates the improvements in terms of complexity and memory usage brought by PEPITA [208] and the Forward-Forward Algorithm, the results show that the Forward-Forward algorithm reduces memory consumption by 40% on average, however it involves additional computation at inference that, can be costly on microcontrollers.
@@ V-E1 Comparison of the usage of the different techniques
We will first start by making a comparison of the different EL techniques’ over the years, by analysing their academic contributions. The figure 4 presents the number of papers for each technique, while the figure 5 references the number of papers (in a logarithmic scale) per technique per year. In order to get the necessary data, we used the advanced search feature of Scopus  [^1] Scopus : https://www.scopus.com/ where we searched the terms ”edge learning”, ”training/learning on the edge/mobile devices”, ”on-device training/learning” and ”on-device adaptation” as well as the relevant keywords for each technique (”Federated Learning”, ”Split Learning”, etc), in the Title, Keywords and Abstract, we made sure to exclude the surveys, conference reviews, books and notes as we are only interested in approaches to train ML models in the edge using the aforementioned techniques. For each family of techniques, we manually reviewed and removed papers that either didn’t provide a contribution, or weren’t about training ML models on edge devices with said technique, despite checking the keywords. Finally, we added a few papers we found manually and either were not available in Scopus or didn’t check the keywords but were relevant in the context of the analysis. The final number of papers associated with the analysis was 511 papers, and some of these papers were briefly covered in V. Note that multiple techniques can be used in a single paper, therefore the total count of techniques in the figures 4 and 5 will exceed 511.
@@
Figure 5: Trend of techniques used to train ML models in the edge over the years
@@ V-E2 Comparison based on the requirements and needs of EL
As discussed in Section IV, several requirements must be met for edge learning, which can function as imprecise measures for assessing the viability of the families of techniques covered in earlier sections. After reevaluating the prerequisites, we eliminated ”Labelled Data Independence” from the comparison, because it is more relevant to the type of machine learning employed and the availability of an autolabeling process than to the approaches under consideration. Additionally, we do not consider ”Task-specific metrics and performance” since that covers multiple metrics used in specific tasks; nevertheless, we introduce a ”High Performance” measure to estimate roughly if the strategies positively or negatively affect the performances. For instance, model compression techniques often lower model performance, whereas incremental and meta-learning typically boost it. Consequently, we employ six distinct measures: ”Computational Efficiency,” ”Memory Footprint,” ”Low Energy Consumption,” ”Quick Training Time,” ”Optimized Bandwidth,” and ”High Performance.” Table II compares the various families of techniques against these requirements.
@@ VI-A Unsupervised Learning
The figure 7 represents the main different types of unsupervised learning approaches used for training ML models in the edge, (a) using unsupervised learning algorithms directly on-device with non-collaborative methods [216, 35]; (b) using unsupervised learning methods to assist in the training of collaborative learning approaches, such as clustered federated learning [79, 217, 218] ; and (c) training an unsupervised learning model on the edge in a collaborative way [219].
@@
Having an unsupervised learning trained on a single edge device is possible if the edge device has enough computation power and/or the learning algorithm is lightweight and can be trained with low resources, the training with such models is no different from the training on the cloud or other devices, the only difference being the constraint of low resources and data available [35]. Examples of unsupervised learning algorithms trained with this approach include [34] that investigates the application of K-Means on mainstream controllers, and [220] that presents the first dedicated CycleGAN accelerator for energy-constrained mobile applications, achieving a higher throughput-to-area ratio and higher energy efficiency than a GPU. In [216], an unsupervised segmentation was proposed that can be executed on edge devices without the need of annotated data. While [221], proposes TMNet, an approach to solve unsupervised video object segmentation problem at the edge. Finally, [222] propose an FPGA based architecture for a Self-Organization Neural Network capable of performing unsupervised learning on input features from a CNN by dynamically growing neurons and connections in order to perform class-incremental lifelong learning for object classification in the edge.
@@
Collaborative learning approaches like FL are promising solution for training ML models in the edge, However, FL the most popular technique on the edge, face multiple challenges such as non-independent and identically collaborative (non-IID) data and uneven computing power [110] and yields suboptimal results when the local clients’ data distributions diverge [79]. To solve these issues, the usage of clustering alongside FL have been proposed multiple times to cluster edge servers with similar environmental data distributions [223]. Among these approaches, Clustered Federated Learning (CFL), introduced by [79] is widely adopted, and it works by exploiting geometric properties of the FL loss surface and grouping the client population into clusters with jointly trainable data distributions. Approaches that build upon CFL have also been proposed, and they include [217, 110, 218, 224]. Other approaches that differ from CFL [79] include [68] that propose a hierarchical over-the-air FL approach which utilizes intermediary servers to form clusters near mobile users; and HPFL-CN [223], a communication-efficient hierarchical personalized federated edge learning framework, that achieves this by utilizing complex network feature clustering to group together edge servers that exhibit similar environmental data distributions. Subsequently, personalized models are trained for each cluster using a hierarchical architecture, resulting in enhanced efficiency. More specifically, HPFL-CN incorporates Privacy-preserving Feature Clustering (PFC) to derive privacy-preserving low-dimensional feature representations for each edge server. This is achieved by mapping the environmental data onto various complex network domains, thereby accurately clustering edge servers with similar characteristics. And Finally, [57] uses unsupervised learning methods on the edge to distinguish between classes across different users, when new classes with completely unseen data distributions emerge on devices in a federated learning setting for audio classification.